OpenStudy (anonymous):

i dont get this Xsquare+5x+6=0

OpenStudy (anonymous):

Factorize the equation first. So you will get (x+3)(x+2)=0

OpenStudy (anonymous):

how did yu do it

OpenStudy (anonymous):

(x+3)(x+2)=0 x+3=0 x+2=0 x=-3 x=-2

OpenStudy (anonymous):

ok so how you went from X^2+5X+6=0

OpenStudy (anonymous):

You need to find to values that when you multiply together, you get 6. so you have only a few choices: 1 * 6 2 * 3 the problem when you use 1*6 is that you cannot get the 5x to be positive.

OpenStudy (anonymous):

Oh sorry. You cannot get the 6 to be positive if you use 1 * 6. Cause your 1 has to become a -1 in order to get the 5x.

OpenStudy (anonymous):

o i see

OpenStudy (anonymous):

actually, we are going to have a trial and error here, We must find factors of 6 that will give us the sum of 5 and that is 2 and 3. so will have (x+3)(x+2)=0 you can check if (x+3)(x+2)=0 is equal to X^2+5X+6=0 by FOIL Method